# Performance testing of a novel short axis photomultiplier tube for the HUNT project

### v1

@@ Abstract
Photomultiplier tubes (PMTs) with large-area cathodes are increasingly being used in cosmic-ray experiments to enhance detection efficiency. The optical modules (OMs) of the High-Energy Underwater Neutrino Telescope (HUNT) have employed a brand new N6205 20-inch microchannel plate photomultiplier tube (MCP-PMT) developed by the North Night Vision Science & Technology (Nanjing) Research Institute Co. Ltd. (NNVT). In order to make the 20-inch PMT fit into the 23-inch diameter pressure-resistant glass sphere, NNVT improved the internal structure of PMT and shortened the height of PMT by more than 10 cm. The first batch of these PMTs has been delivered for preliminary research work. This paper describes a specific PMT testing platform built for the first batch of 15 MCP-PMTs, and some performance parameters of PMT, such as P/V ratio, TTS and nonliniearity, are measured.The measurement results show that the new PMT still has good performance and can meet the requirements of HUNT project.
@@ 3.1 Dark room
In the laboratory, a separate room was set up as a darkroom, where two barrels were installed for testing the PMTs. The walls of the darkroom and the outer wall of each barrel were pasted with materials for shielding the magnetic field, resulting in a residual magnetic field at the PMT cathode position of less than 5 $\mu$T. The magnetic shielding material is black, and a black cloth is pasted inside each barrel to eliminate the influence of stray light. The size of a barrel is shown in figure 4. It has an inner diameter of 540 mm and a height of 900 mm. Three 135° brackets were evenly installed on the inner wall of each barrel at a height of 670 mm for PMT placement. For these setting, the top of the PMT cathode is 500 mm above the base, and the PMT equator is 670 mm above the base.
@@ 4.1 Waveform inspection
Figure 6: The waveform under SPE conditions after amplified tenfold, showing an amplitude of 60 mV.
@@ 4.2 Cooling
When cooling tests are initially conducted, it is essential to ensure that the tests are performed under SPE conditions. Therefore, before running the test script, a debugging script is executed to determine the HV and trigger ratio during testing. During debugging, the high voltage is set initially according to the values provided in the NNVT factory test table, which calls for taking a 500,000 trigger signal, and repeating four times. The output includes the statistical average and error of the trigger ratio, gain, and charge. The high voltage and other settings are adjusted to achieve trigger ratios and gains at levels of 10% and 5×106, respectively, before running the formal test script.
The cooling test includes monitoring four parameters: trigger ratio, gain, charge, and DNR. This serves as an evaluation data set and is conducted over a period of at least 2 hours. The signal output from the PMT is routed through a switch, with the switch set to the first gear. The combined signal is then amplified tenfold using an N979 amplifier before being processed by the V792N.
@@
The instability during testing comes from the environment, leakage current, residual gases in the PMT, etc. Therefore, consistent working conditions need to be ensured during the testing process. The cooling tests are considered complete when the trigger ratio variation is less than 2%, gain variation is less than 3%, charge variation is less than 2%, and DNR variation is less than 10%.
@@ 4.5 Q-T effect
A 2-kHz signal trigger rate is set, and 10,000 trigger signals are collected for both QDC and TDC. The light transmission rate of the attenuator is varied, covering a total of 30 points, and at each point, measurements of both time and charge are made. The amount of charge and time for each transmittance point are calculated, with a time error of $\text{FWHM}/{\sqrt{12}}$. The FWHM is obtained by fitting the time distribution with a Gaussian, as shown in the figure 10.
@@ 4.7 Afterpulse ratio
The afterpulse ratio test measures the slow afterpulse components in the PMT signal. After completing the anode nonlinearity test, the light intensity is adjusted to the 30 PEs level. The PMT signal is then input into the N979 and amplified tenfold. It is next input into the CIQTEK TDC0110. This TDC module can adjust the threshold through PC driver software, with the threshold set to 1/3 PE. The plotting window is set to 10 $\mu$s, and the acquisition time is set to 4 minutes. Clicking the start button to begins the test; the results are shown in figure 12. The afterpulse ratio is calculated using equation (4.5). The count from 0.5 $\mu$s to 10 $\mu$s after the main pulse is determined statistically and called $N_{A}$. Under the same conditions, DNR measured during the same time window when the light source is turned off is $D_{R}$$\times$ 9500 $\times$ 10-9. $N_{0}$ is the count of the main pulse. The afterpulse ratio, given by equation (4.5) is 0.86%.
@@
Figure 12: Time distribution of the afterpulse.
@@ 5 Summary
We set up a set of PMT test platform and established the test process of PMT performance. The test results of HV, peak-to-valley ratio, DNR, TTS, and SPE resolution and Afterpulse ratio are summrized in figure 13. These parameters will be used to evaluate the performance of PMTs for the HUNT experiment. All PMTs operate at a gain of 5 $\times$ 106; the high voltages range vary from 1650 to 1800 V, peak-to-valley ratios all are larger than 2.0, and SPE resolutions are smaller than 50%. After a certain cooling period, the DNRs become below 25 kHz. And TTSs’ values are less than 7 ns. Afterpulse ratios of the tested PMT are all below 0.1%. Therefore ,all above test results have met the requirements.

### v2

@@ Abstract
Photomultiplier tubes (PMTs) with large-area cathodes are increasingly being used in cosmic-ray experiments to enhance detection efficiency. The optical modules (OMs) of the High-Energy Underwater Neutrino Telescope (HUNT) have employed a brand new N6205 20-inch microchannel plate photomultiplier tube (MCP-PMT) developed by the North Night Vision Science & Technology (Nanjing) Research Institute Co. Ltd. (NNVT). In order to make the 20-inch PMT fit into the 23-inch diameter pressure-resistant glass sphere, NNVT improved the internal structure of PMT and shortened the height of PMT by more than 10 cm. The first batch of these PMTs has been delivered for preliminary research work. This paper describes a specific PMT testing platform built for the first batch of 15 MCP-PMTs, and some performance parameters of PMT, such as peak-to-valley ratio, TTS and nonliniearity, are measured. The measurement results show that the new PMT still has good performance and can meet the requirements of HUNT project.
@@ 3.1 Dark room
In the laboratory, a separate room was set up as a darkroom, where two barrels were installed for testing the PMTs. The walls of the darkroom and the outer wall of each barrel were pasted with iron-based amorphous alloy foils [9] for shielding the magnetic field, resulting in a residual magnetic field at the PMT cathode position of less than 5 $\mu$T. The magnetic shielding material is black, and a black cloth is pasted inside each barrel to eliminate the influence of stray light. The size of a barrel is shown in figure 4. It has an inner diameter of 540 mm and a height of 900 mm. Three 135° brackets were evenly installed on the inner wall of each barrel at a height of 670 mm for PMT placement. For these setting, the top of the PMT cathode is 500 mm above the base, and the PMT equator is 670 mm above the base.
@@ 4.1 Waveform inspection
Figure 6: The waveform under SPE conditions after being amplified by a factor of ten, showing an amplitude of 60 mV.
@@ 4.2 Cooling
[…]
The cooling test includes monitoring four parameters: trigger ratio, gain, charge, and DNR. This serves as an evaluation data set and is conducted over a period of at least 3 hours. The signal output from the PMT is routed through a switch, with the switch set to the first gear. The combined signal is then amplified tenfold using an N979 amplifier before being processed by the V792N.
@@
The instability during testing comes from the environmental temperature, leakage current, residual gases in the PMT [13]. Therefore, consistent working conditions need to be ensured during the testing process. The cooling tests are considered complete when the trigger ratio variation is less than 2%, gain variation is less than 3%, charge variation is less than 2%, and DNR variation is less than 10%.
@@ 4.5 Q-T effect
A 2-kHz signal trigger rate is set, and 10,000 trigger signals are collected for both QDC and TDC. The light transmission rate of the attenuator is varied, covering a total of 30 points, and at each point, measurements of both time and charge are made. The amount of charge and time for each transmittance point are calculated, with a time error of $\text{FWHM}/{\sqrt{12}}$. The FWHM is obtained by fitting the time distribution with a Gaussian. The test results of Q-T effect are shown in figure 10.
@@ 4.7 Afterpulse ratio
Afterpulse is a signal generated by the ionization of trace gases remaining in the PMT by photoelectrons, which return to the photocathode and are typically microsecond slower than the main pulse [14]. The afterpulse ratio test measures the slow afterpulse components in the PMT signal. According to Section 4.1, SPE corresponds to 6 mV, and the signal amplitude displayed by the PMT on the oscilloscope is calibrated to 180 mV, that is, 30 PE, by adjusting the light intensity through an optical attenuator. This is because in the SPE state, the PMT signal-to-noise ratio of PMT is relatively low, resulting in significant test errors. According to the conclusion in Section 4.6, the level of 30 PEs does not cause any non-linearity in the test results. The PMT signal is then input into the N979 and amplified tenfold. It is next input into the CIQTEK TDC0110. This TDC module can adjust the threshold through PC driver software, with the threshold set to 1/3 PE. The plotting window is set to 10 $\mu$s, and the acquisition time is set to 4 minutes. Clicking the start button to begins the test; the results are shown in figure 12.
The afterpulse ratio is calculated using equation (4.5). The count from 0.5 $\mu$s to 10 $\mu$s after the main pulse is determined statistically and called $N_{A}$. Under the same conditions, DNR measured during the same time window when the light source is turned off is $D_{R}$$\times$ 9500 $\times$ 10-9. $N_{0}$ is the count of the main pulse. The denominator of 30 is used to normalize the result to SPE. The afterpulse ratio given by equation (4.5), taking PAB2308-9022 as an example, is 0.086% $\pm$ 0.001%.
@@
Figure 12: Time distribution of PAB2308-9022 afterpulse. On the right of the black dotted line in the figure is the part of the afterpulse.
@@ 5 Summary
We set up a set of PMT test platform and established the test process of PMT performance. The test results of HV, peak-to-valley ratio, DNR, TTS, and SPE resolution and Afterpulse ratio are summrized in figure 13. These parameters will be used to evaluate the performance of PMTs for the HUNT experiment. All PMTs operate at a gain of 5 $\times$ 106; the high voltages range vary from 1650 to 1800 V, peak-to-valley ratios all are larger than 2.0, and SPE resolutions are smaller than 50%. After a certain cooling period, the DNRs become below 25 kHz. And TTS are less than 7 ns. Afterpulse ratios of the tested PMT are all below 0.1%. Table 3 gives the statistical analysis of the 15 PMT test results mentioned above. Therefore, all above test results have met the requirements.
@@
Table 3: The mean and standard deviation of the measured 15 PMTs.
Parameter Value
HV (V) -1743 $\pm$ 28
Peak-to-valley 2.96 $\pm$ 0.28
DNR (kHz) 13.13 $\pm$ 3.65
TTS (ns) 6.22 $\pm$ 0.25
SPE Resolution (%) 40.37 $\pm$ 2.61
Afterpulse ratio (%) 0.06 $\pm$ 0.02